OpenStudy (madgirlwithabluebox):

Simplify the expression. (–3i)( –4i)

OpenStudy (madgirlwithabluebox):

@UnkleRhaukus

OpenStudy (unklerhaukus):

\[(-3i)(-4i)\\=(-1\times3\times i)\times(-1\times4\times i)\]we can rearrange \[=(-1\times-1)\times(3\times4)\times(i\times i)\\=\]

OpenStudy (madgirlwithabluebox):

1 x 12 x i^2?

OpenStudy (madgirlwithabluebox):

12i^2?

OpenStudy (unklerhaukus):

\[=1\times12\times i^2\\=12i^2\]good. Do you know what \(i^2 =?\)

OpenStudy (madgirlwithabluebox):

no?

OpenStudy (unklerhaukus):

\[\boxed{i = \sqrt{-1}}\qquad\implies\qquad i^2=\sqrt{-1}\times\sqrt{-1}=-1\]

OpenStudy (madgirlwithabluebox):

so i^2=-1

OpenStudy (unklerhaukus):

yeah, which makes \[=12i^2\\=12(-1)\\=\]

OpenStudy (madgirlwithabluebox):

-12

OpenStudy (unklerhaukus):

Yes!

OpenStudy (unklerhaukus):

But if you haven't been studying complex numbers, maybe \(i\) is not supposed to be the imaginary unit, in which case you should just leave your answser as \(12i^2\)

OpenStudy (madgirlwithabluebox):

ok.. i have 1 last question. someone gave me the answer but i wanted to how how to do it. could you help me one last time?

OpenStudy (unklerhaukus):

ok!

OpenStudy (madgirlwithabluebox):

than you so so so much!